OpenStudy (anonymous):

Calculate the amount of your simple interest investment after 10 years. Remember to use the same rate and compounding as in #1. So my interesrt is .84% daily. and i already have 7500 in my account

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large I=Prt\] Hmm I don't think simple interest has anything to do with compounding as we had in our compound interest problem. I wonder why they mentioned it here. hmm :o

zepdrix (zepdrix):

\[\Large P=7500\]\[\Large r=.0084\]\[\Large t=10\] \[\Large I=?\] :D

OpenStudy (anonymous):

630?

zepdrix (zepdrix):

Mmmmmmm yah that sounds right :o I'm not too familiar with the interest problems but I think that's all they wanted us to do for this one.

OpenStudy (anonymous):

Thank you so much for helping me! yayy!

zepdrix (zepdrix):

lol yer silly c:

OpenStudy (anonymous):

OK ARE YOU UP FOR ONE MORE?"_

zepdrix (zepdrix):

sure, just post it in this thread, it's not very long yet so shouldn't be a problem.

OpenStudy (anonymous):

let me just send you a picture hold on

OpenStudy (anonymous):

ok we are on number 3 :))))

zepdrix (zepdrix):

Ooo fun graph tool :O look at all the buttons!

OpenStudy (anonymous):

ahahhahhaaahah! i just laughed so hard hahahaha your so silly!

zepdrix (zepdrix):

sure, just post it in this thread, it's not very long yet so shouldn't be a problem.

zepdrix (zepdrix):

Ooo fun graph tool :O look at all the buttons!

zepdrix (zepdrix):

So let's graphing the Simple Interest first since that will be easier. We'll graph `time` in the horizontal direction (the x direction), and `cash monies` in the vertical direction (the y direction).

zepdrix (zepdrix):

So this is where we'll want to go back to using a ton of brackets so our answer will be ACCURATE.

zepdrix (zepdrix):

Recall after we plugged in all of our jimmy jammies we had,\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365*10}\] But what we want to do now is, we DON'T want to plug in the `t` value. We want to allow our `t` to vary.Recall after we plugged in all of our jimmy jammies we had,\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365*t}\]

zepdrix (zepdrix):

Blah fail copy paste.. my bad.

zepdrix (zepdrix):

So that function we have there... we want to write it using brackets and stuff so the program will graph it.

OpenStudy (anonymous):

woah woah woah what? i have no idea what your talking about :(

zepdrix (zepdrix):

You remember this?\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365*10}\]

OpenStudy (anonymous):

and here is the link to the graph thing. http://www.geogebra.org/cms/en/

zepdrix (zepdrix):

There's a chrome app? oo very cool

OpenStudy (anonymous):

and yes i do remember that equation

zepdrix (zepdrix):

\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365*\color{#F35633}{10}}\] The 10 is there because the problem told us to \(\Large t=10\). We want to undo plugging that 10 in giving us this now.\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365*\color{#F35633}{t}}\]k? :o

OpenStudy (anonymous):

ok! and what does this have to do with the graphing though???

zepdrix (zepdrix):

The amount of cash money we can make, \(\Large A\) is related to the amount of time, \(\Large t\), we allow to pass. So we can show how our money relates to time by graphing a line, or curve in this case.

zepdrix (zepdrix):

Ah crap, I said we were going to start with `Simple Interest` didn't I? Grr

zepdrix (zepdrix):

Ok ok let's switch to the Simple Interest Formula a sec.

zepdrix (zepdrix):

\[\Large I=Prt \qquad\to\qquad I=7500*.0084*\color{#F35633}{10}\]Again, we want to undo plugging in the 10.\[\Large I=7500*.0084*\color{#F35633}{t}\]

zepdrix (zepdrix):

The graphing utility will only allow us to graph in x and y. So remember what I mentioned earlier. We'll graph time in the horizontal direction (as x), and Interest in the vertical direction (as y).

zepdrix (zepdrix):

\[\Large y=7500*.0084*x\] Understand? D: A little bit maybe?

zepdrix (zepdrix):

Uh oh :c your head esplode?

OpenStudy (anonymous):

no no no i understand!

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

So at the bottom of the graphing utility, there is a space to type in your function.

zepdrix (zepdrix):

Did you try graphing it yet? :o

OpenStudy (anonymous):

what do i graph?

OpenStudy (anonymous):

ugh this is so frustrating!

zepdrix (zepdrix):

This is our simplie interest formula, graph this, \[\Large y=7500*.0084*x\]

OpenStudy (anonymous):

in x?

zepdrix (zepdrix):

I changed \(\Large I\) to \(\Large y\), and \(\Large t\) to \(\Large x\). That allows us to graph this in \(\large x\) and \(\Large y\).

OpenStudy (anonymous):

so 7500 * . 0084 * x is for the Y axis?

zepdrix (zepdrix):

yes, x is for the x axis. 7500*.0084*x is for the y axis.

zepdrix (zepdrix):

You see the `input` box at the bottom, right? :o

OpenStudy (anonymous):

yes and i entered it in there

OpenStudy (anonymous):

and this is what it looks like

OpenStudy (anonymous):

is this correct?

zepdrix (zepdrix):

Yes. So you see that tiny sliver? It's a line with a really really big slope. So it's growing really quickly.

zepdrix (zepdrix):

The other one will be a tad harder to type into our input box. We'll have to be careful with the brackets we use.\[\Large A=7500\left(1+\frac{.0084}{365}\right)^{365t}\] In x and y this becomes,\[\Large y=7500\left(1+\frac{.0084}{365}\right)^{365x}\]

zepdrix (zepdrix):

We'll want to input it like this,\[\Large y=7500(1+(.0084/365))\wedge(365x)\]

zepdrix (zepdrix):

Hmm it's not accepting it.. lemme see what's going on -_-

OpenStudy (anonymous):

ok:(

zepdrix (zepdrix):

Mmm ok I guess we don't want to use that many brackets. put this into the input box: 7500(1+.0084/365)^(365x)

zepdrix (zepdrix):

You'll have to zoom way way out in order to actually see the line.

OpenStudy (anonymous):

ok lets seeee....

OpenStudy (anonymous):

how do i zoom?

zepdrix (zepdrix):

Do you have a scroll wheel on your mouse? :) Are you using a mouse? hehe

OpenStudy (anonymous):

no im on a laptop

zepdrix (zepdrix):

Go to the last icon on the top bar, which looks like a cross with arrows on the end. Click the white arrow under it for a drop down menu, there should be a zoom out option.

OpenStudy (anonymous):

even when i zoom in i cant see it :(

zepdrix (zepdrix):

zoom out! :O

OpenStudy (anonymous):

does your look like this?

zepdrix (zepdrix):

yesss good good

zepdrix (zepdrix):

So hopefully we're doing this right +_+ lol

zepdrix (zepdrix):

Mmm crap hold on....

OpenStudy (anonymous):

i know right haha now what?

zepdrix (zepdrix):

Ohhh Ok ok ok ok ok ok ok ok

zepdrix (zepdrix):

Check back on your first problem page. See below the input example, the message: `Notice that A= was not included, that 10 was used instead of 10,000 (because this graph is scaled in thousands),`

zepdrix (zepdrix):

I think they want us to scale this like they did in the example. Hmm

zepdrix (zepdrix):

It will look a lot nicer when we graph it.

OpenStudy (anonymous):

how do we do that?

zepdrix (zepdrix):

Ummmmmmmmmm

OpenStudy (anonymous):

hahah i thought so too!

zepdrix (zepdrix):

I guess we could use the same scale they did in the example. So we would divide our `Principle` value by 10,000. Err no, 1000 scale would make more sense maybe.... grrrr

OpenStudy (anonymous):

what should i do then??

zepdrix (zepdrix):

Dividing our principal value by 1000 gives us, P=7500/1000 --> P=7.5 So we would instead want to graph: 7.5(1+.0084/365)^(365x) and 7.5*.0084x

zepdrix (zepdrix):

They look a little bit better I guess. I dunno.. the instructions are confusing for this question >:C

zepdrix (zepdrix):

You'll have to zoom OUT a bit to get the nice shape. It will look better than the previous one did though.

OpenStudy (anonymous):

i got it!but how come it doesnt look like the one in the example?

zepdrix (zepdrix):

What example ? :o

OpenStudy (anonymous):

the one that i linked to you? ill show it again hold on

zepdrix (zepdrix):

There is no line graphed in that picture.. You mean the way the graph layout looks?

OpenStudy (anonymous):

yes! idk is it suppossed to look like the example?

zepdrix (zepdrix):

Mmm there was a button to add a `grid` .. i'm trying to remember where it was at.

zepdrix (zepdrix):

I guess you can right click on the graph, `show grid`

zepdrix (zepdrix):

That will make it look a bit similar :o

OpenStudy (anonymous):

where is that at?

zepdrix (zepdrix):

right click your mousey button on the graph. a menu will pop up

OpenStudy (anonymous):

oh ok! now what do i do?

zepdrix (zepdrix):

You can label the axes like they did in the example, if you want. Something like this, but not stupid lol.

zepdrix (zepdrix):

That's probably not necessary though :o lol

zepdrix (zepdrix):

you are online school or something?